# MathGAP: Out-of-Distribution Evaluation on Problems with Arbitrarily Complex Proofs

### v1

@@ 3.2 Chain-of-thought solutions as proof trees
Solving math word problems is a form of deductive reasoning, where the correct answer follows necessarily from what is stated in the text through a combination of rules of arithmetic and world knowledge. We use such derivations, called proofs, to characterize the structure of problems. The proofs in our setup are analogous to those found in other deductive systems, e.g., of first-order logic.author=andreas,color=blue!40,size=,fancyline,caption=,]This could be a good place to cite other NLP papers which have formalized reasoning through deduction systems. Any suggestions? abu:there is a lot we could cite here; maybe going all the way back to Montague, surveys of formal semantics, categorial grammar (where parsing itself is formalized as a reasoning problem)
@@
A math word problem can be solved by applying inference rules until the answer of the problem has been proved. Formally, a proof tree is a rooted ordered tree where each node is labeled with a logical form $L$, which is the conclusion of a proof step with premises $L_{1},\dots,L_{N}$, matching the labels of the child nodes in the same order. Leaf nodes are thus labeled with axioms. A proof of a particular problem is a proof tree where the labels on the leaf nodes are logical forms in the body of the problem, and the label at the root is the answer to the problem. The tree shown in Fig. 1 illustrates a proof of the problem given in the same figure.
Proof complexity.
We seek a systematic way to analyze whether LLMs can generalize from simple to more complex proofs. The definition of proof given above gives rise to several ways of characterizing reasoning, out of which we consider four: linearity, depth, width, and ordering. These are defined below.
@@
We say that a proof tree is linear if every one of its inference rules takes at most one premise that is not an axiom, and nonlinear otherwise. In other words, every step in the solution to a linear problem only uses at most one fact not directly present in the problem. The proof tree given in Fig. 1 is nonlinear because the proof step that deduces the answer uses two premises that are not themselves axioms.author=andreas,color=blue!40,size=,fancyline,caption=,]We should give an example of a linear problem as well so that this is easier to understand abu: agreed, maybe in appendix? The depth of a proof tree is defined as its height, i.e., the maximum path length from the root node to any leaf node. Next, the width of a proof tree is the number of leaf nodes, or axioms, that it contains. The proof tree in Fig. 1 thus has depth 2 and width 5. In words, this means that the answer is two proof steps “away” from any of the axioms, and that the problem has five (declarative) sentences. [^4] In this paper, the width of the proof tree matches the number of declarative sentences (i.e., all sentences except the question) in the corresponding problem since we restrict the logical forms and sentences to be one-to-one. In general, that need not be the case since a single sentence could be written as a conjunction of several logical forms. One could also imagine problems that have irrelevant sentences, which do not represent any logical form at all. Lastly, we define the canonical ordering to be the ordering of leaf nodes as visited left-to-right. Note that the sentences in the natural language annotation from Fig. 1 follows the canonical ordering of the proof tree. Informally, this is the easy way of ordering the sentences in the problem, since it matches the ordering of the proof steps.author=andreas,color=blue!40,size=,fancyline,caption=,]Would be great if we could find support for this statement in cogsci literature However, there exist multiple alternative orderings in general. In particular, given a proof tree that only contains commutative inference rules (see footnote 3), all orderings of the leaf nodes are valid. For instance, swapping the ordering of the first two leaf nodes in the tree in Fig. 1 yields a problem with the same proof.
@@ 4 Evaluation Framework
Equipped with the formalism explained in § 3, we propose an evaluation framework which tests LLMs on MWPs of arbitrary proof complexity, called MathGAP. In § 4.1, we explain the generation method behind MathGAP. We then outline how this method can be used for OOD evaluation in § 4.2.
@@ 4.1 Generation Method
We propose a method for synthetic generation of problems and their CoT solution explanations. It consists of three high-level steps: (i) sample a proof tree, (ii) map the logical forms at the leaf nodes into a natural language problem, and (iii) map the proof steps into a CoT solution. [^5] Jin et al. (2023) presents a related method for generating causal reasoning problems. Fig. 1 gives an illustration of these steps; note that the agent properties are abbreviated for brevity.
@@
In detail, the procedure is as follows: We first sample a logical form at the root of the proof tree. We then sample a specific inference rule as well as its premises. We recursively repeat this procedure for each premise until a predetermined stopping criterion has been reached, resulting in a proof tree. The stopping criterion is user-specified; it could be, e.g., when the tree has reached a certain depth or width. The categorical properties for the logical forms (i.e., agent, entity, attribute, and unit) are sampled uniformly at random from handwritten vocabularies. The quantities are sampled uniformly at random from a user-specified range; in our study presented in § 5 we use $2{-}20$.
@@
We then traverse the leaves of the proof tree in some order and convert each into a sentence via a natural language template, forming the text corresponding to the body of the problem. The default order follows the proof tree’s canonical ordering, but other orders are possible as well. The sentences are sampled uniformly at random from a predefined set of templates; some example templates are given in Table 1. The question is converted from the logical form at the root of the proof tree, and is sampled from a set of interrogative sentences.
@@
In the final step, we generate the CoT annotation for the problem. We first obtain a sequential ordering of the CoT proof steps by visiting the nodes in the proof tree with a post-order traversal. [^6] Note that the choice of traversing the nodes in post order corresponds to applying each inference rule as soon as all of its premises are available. In addition, post-order traversal visits the axioms in the canonical ordering. Alternative orders of tree traversal would violate at least one of these constraints.author=andreas,color=blue!40,size=,fancyline,caption=,]If we want to comment on similarity to the ordering of human reasoning etc we should look into literature on that. Start with the figural effect reference by Eisape et al. (2024) and the proof ordering references by Boguraev et al. (2024) We then convert each proof step into a sentence via a natural language template. The templates used for the axioms are identical to those used in the body of the problem text. In other words, the proof steps corresponding to axioms simply repeat the relevant part of the problem text. For the other proof steps, we have templates which are designed to explain the computation step that is performed. Fig. 1 illustrates the mapping between nodes in the proof tree and sentences in the CoT annotation. For instance, the third sentence in Fig. 1 explains the conclusion cont(Lucy, 27, apple), which is deduced from the two leftmost nodes in the tree: “So Lucy has $17+10=27$ apples”.
Our pipeline is similar to the one proposed by Opedal et al. (2024), except that we use a recursive procedure to generate each proof step starting at the root, instead of a sequential one. This is a more general approach which also enables generation of nonlinear proofs. [^7] The comp-eq rule enables the generation of nonlinear proof trees with arbitrary depth and is not included in the list of relations from Opedal et al. (2023).author=andreas,color=blue!40,size=,fancyline,caption=,]We need to explain the nuance of building nonlinear proofs in an appendix. We can create a nonlinear ones of depth two with only comp , but for deeper than that we need comp-eq. In the experiments we sample depth 2 examples of both comp and comp-eq.
@@ 4.2 MathGAP: Mathematical Generalization on Arithmetic Proofs
With the generation method just described, we can generate new synthetic problems where we control for the structure of the proof and its complexity. MathGAP can be used for various forms of systematic studies on arithmetic reasoning; here, we focus on OOD generalization. Specifically, we can select a family of problems of interest (e.g., linear problems under comp inference rules) and generate a train-test split of problems for which the problems in the test set are more complex than those in the training set (e.g., containing problems with deeper proofs).
The generation method can be flexibly adapted to fit specific studies. One can add new logical forms and inference rules, restrict the choice of existing ones during sampling, or create logical forms for irrelevant sentences (Shi et al., 2023). One can also vary the text templates for the logical forms and inference rules, the vocabularies for agents, entities, attributes and units, and the range of numbers from which the quantities are sampled. In addition, one may increase linguistic diversityauthor=andreas,color=blue!40,size=,fancyline,caption=,]There must be a better, more linguistically correct term for this by incorporating a last step in which a language model paraphrases the templated texts, or, alternatively, use a procedure like the one proposed by Mishra et al. (2024). In our study we do not perform such paraphrasing, since we want to guarantee that the sentences are faithful to the proof trees from which they are generated. However, future studies using MathGAP may choose to prioritize differently.
5 Experiments
author=andreas,color=blue!40,size=,fancyline,caption=,]Remind the reader somewhere here of why our evaluation framework is neat. We could have done an analysis on, say, GSM8k, but we prefer our synthetic data so that we can (i) avoid data contamination and (ii) control the proof tree characteristics
We apply MathGAP to study how LLMs generalize from proof demonstrations of simple problems in context to solve problems of higher complexity at inference. We test OOD generalization through several sets of experiments: depth generalization for linear problems (§ 5.1), width generalization for linear problems (§ 5.2), depth generalization for nonlinear problems (§ 5.3), and generalization to permutations (§ 5.4). We are also interested in how the choice of the distribution of problems provided in context affects performance, which is often not obvious (Min et al., 2022; Turpin et al., 2023).
@@ 5 Experiments
For each set of experiments, the general experimental setup is as follows: We generate five test sets of different degrees of complexity with 400 problems in each. We then generate model predictions for the problems in these test sets under four different prompts: (i) A zero-shot baseline prompt, which contains only the test problem; (ii) A prompt of primitive examples, in which all problems contain only one proof step of the same inference rule that is used in the test problem; (iii) A prompt of examples within a range of complexities, all of which are simpler than the test problem;author=abu,color=orange!40,size=,fancyline,caption=,]maybe we should format these settings differently so that it’s clearer that we will refer to them later in this section; either by bolding or maybe by naming them “Prompt (i):” (iv) An in-distribution baseline prompt, containing examples that are of the same complexity as the test problem.
[…]
Responses are generated using greedy decoding and a maximum length of 4,096 tokens. Model predictions are obtained by extracting the last number occurring in the model’s generated output. We report answer accuracies and compute 95% confidence intervals using bootstrap sampling (Efron, 1992). We evaluate the following models on all test sets: Mixtral-8x7B (Jiang et al., 2024), Llama3 with 8B and 70B parameters (Llama Team, 2024), GPT-3.5 Turbo and GPT-4o (OpenAI, 2024). [^10] The maximum context windows for these models are 65k, 8k, 8k, 16k, 128k, respectively. Combined with a maximum number of output tokens of 4,096, the models have a theoretically sufficient context size to solve all test problems. The largest test problems (nonlinear depth 6) have around 2.5k tokens, including CoT annotation. For the OpenAI models we used “gpt-3.5-turbo-0125” and “gpt-4o-2024-05-13” as the model ids for all experiments.
Figure 2: Answer accuracies for generalization to increasing depth and width for linear problems, across models and in-context distributions. Depth is increased using inference rules involving comp and transfer . Width is increased using partwhole.
@@ 5.1 Linear Depth Generalization
We consider linear problems using both comp and transfer .author=andreas,color=blue!40,size=,fancyline,caption=,]We should give example proof trees for all our settings in appendix The five test sets consist of problems with depths between $6{-}10$. As such, the examples provided in the range of setting (iii) described above have depths between $1{-}5$. We give an example of a problem with depth $6$ in § B.1.
@@ Results.
We illustrate the results in the top and middle rows of Fig. 2. As expected, we observe a decreasing trend in performance as depth increases, for all in-context distributions. Curiously, GPT-3.5-Turbo is the only model for which adding in-context examples consistently leads to improvements. For the majority of cases, performance tends to be larger when providing a range of examples of varying complexity as compared to only primitive examples. The ranged OOD cases often show comparable performance to that of in-distribution context, suggesting that these models benefit from demonstrations on how problems of varying complexity are solved. Llama3-8B is an outlier in that its solving accuracy for in-distribution context is (substantially) lower as compared to other prompts. When inspecting the model outputs, we see that the model often fails to answer the inference question, recognizing instead that it has been provided a sequence of in-context examples. This could potentially be a consequence of how Llama3 is trained, using smaller context windows for the initial stages of pretraining (Llama Team, 2024). Finally, we note that GPT-4o and the larger Llama3 model show close to perfect performance across all test sets.
@@ 5.2 Linear Width Generalization
To test generalization to greater proof width, we use partwhole problems with a fixed depth of $1$. The five test sets have widths between $6{-}10$ and the width of the problems in setting (iii) described above have widths between $1{-}5$.
@@ Results.
The results are shown in the bottom row of Fig. 2. For GPT-3.5 and Mixtral-8x7B, we observe that providing in-context examples improves accuracy over zero-shot in most cases. Like for linear depth generalization, it seems that having a range of OOD examples of varying complexity is superior to only including primitive examples. Llama3-8B exhibits worse performance with in-distribution context here as well. In general for the Llama models, adding OOD in-context examples usually leads to lower performance over zero-shot. Apart from GPT-3.5-Turbo, the performance is generally lower as compared to a deep problem with the same number of sentences (§ 5.1), often with a sharper decreasing trend. Such is the case even for Llama3 70B, which performed almost perfectly in the linear depth setting. This would suggest that it is more difficult to generalize in terms of width than in terms of depth. Again, GPT-4o demonstrates near perfect performance across all test sets.
Figure 3: Answer accuracies for generalization to increasing depth and width for nonlinear problems, across models and in-context distributions. Depth is increased using inference rules involving comp and comp-eq . The in-distribution contexts were too large to fit in the prompts for the Llama3 models and GPT-3.5 Turbo at some of the higher depths. App. C shows additional results on o1-preview.

### v3

@@ 3.2 Chain-of-thought solutions as proof trees
Solving MWPs is a form of deductive reasoning, where the correct answer follows from what is stated in the text through a combination of rules of arithmetic and world knowledge. We use such derivations, called proofs, to characterize these problems. The proofs in our setup are analogous to those in other proof systems, such as natural deduction Gentzen (1935); Pfenning (2004). [^3] In fact, each of our inference rules shown in § 3.1 can be decomposed into the primitive inference rules of natural deduction, using first-order logic expressions. App. B demonstrates by example how this is done.
@@
A math word problem can be solved by applying inference rules until the answer to the problem has been proved. Formally, a proof tree is a rooted ordered tree where each node is labeled with a unique logical form $L$, which is the conclusion of a proof step with premises $L_{1},\dots,L_{N}$, matching the labels of the child nodes in the same order. Leaf nodes are thus labeled with axioms. A proof of a particular problem is a proof tree where the labels on the leaf nodes are the logical forms in the body of the problem, and the label at the root is the answer to the problem. The tree shown in Fig. 1 illustrates a proof of the problem given in the same figure.
Complexity of proof trees.
@@
We seek a systematic way to analyze whether LLMs can generalize from simple to more complex proofs. [^5] Note that our use of the term “complexity” is different from its conventional use in proof theory, in which proof complexity refers to the efficiency of a proof system (Cook & Reckhow, 1979). The definition of proof given above gives rise to several ways of characterizing reasoning, out of which we consider four: depth, width, shape, and ordering. The depth of a proof tree is defined as its height, i.e. the maximum path length from the root node to any leaf node. Next, the width of a proof tree is the number of leaf nodes, or axioms, that it contains. The proof tree in Fig. 1 thus has depth 2 and width 5. In words, this means that the answer is two proof steps “away” from any of the axioms and that the problem has five (declarative) sentences. [^6] In this paper, the width of the proof tree matches the number of declarative sentences (i.e., all sentences except the question) in the corresponding problem since we restrict the logical forms and sentences to be one-to-one. In general, this need not be the case since a single sentence could be written as a conjunction of several logical forms. One could also construct problems that have irrelevant sentences which either represent logical forms that will not be used in the proof or have no corresponding logical form at all. We distinguish two forms of proof-tree shapes: We say that a proof tree is linear if each of its proof steps takes at most one premise that is not an axiom, and nonlinear otherwise. [^7] We note that our notion of linear proofs is distinct from linear logic (Girard, 1987). In other words, every step in the solution to a linear problem uses only at most one fact not directly present in the problem; see Fig. 6 for an example. The proof tree given in Fig. 1 is nonlinear because the proof step that deduces the answer uses two premises that are not themselves axioms. Lastly, we define canonical ordering as the ordering of the leaf nodes when visited from left to right. Note that the sentences in the natural language annotation from Fig. 1 follow the canonical ordering of the proof tree. Informally, this is an easy way to order the sentences in the problem, since it matches the ordering of the proof steps. However, there exist multiple alternative orderings in general. In particular, given a proof tree that only contains commutative inference rules (see footnote 4), all orderings of the leaf nodes are valid. For instance, swapping the ordering of the first two leaf nodes in the tree in Fig. 1 yields a problem with the same proof.
@@ 4 Evaluation Framework
Equipped with the formalism explained in § 3, we propose an evaluation framework that tests LLMs on MWPs that have arbitrarily complex proofs, called MathGAP. In § 4.1, we explain the generation method behind MathGAP. We then outline how this method can be used for OOD evaluation in § 4.2.
@@ 4.1 Generation Method
We propose a method for synthetic generation of problems and their CoT solution explanations. The approach consists of three high-level steps: (i) sample a proof tree, (ii) map the logical forms at the leaf nodes into a natural language problem, and (iii) map the proof steps into a CoT solution. [^8] Our pipeline is similar to Opedal et al.’s (2024), except that we use a top-down procedure to generate full proof trees, rather than a sequential one to generate only axioms. This is a more general approach that enables generation of proofs with arbitrary shapes, e.g., nonlinear ones using comp-eq ; see App. D. Our pipeline is also similar to the method proposed by Jin et al. (2023) for generating causal reasoning problems. Fig. 1 gives an illustration of these steps; note that the agent properties are abbreviated for brevity.
@@
In detail, the procedure is as follows: We first sample a logical form to label the root of the proof tree. We then sample an inference rule whose conclusion matches the root, and use it to connect the root node to child nodes representing the premises of the sampled inference rule. We repeat this procedure recursively for each premise at the leaf nodes of the current tree until a predetermined stopping criterion has been reached, resulting in a proof tree. The stopping criterion is user-specified; it could be, e.g., when the tree has reached a certain depth or width. The categorical properties for the logical forms (i.e., agent, entity, attribute, and unit) are sampled uniformly at random from some vocabularies.
@@
We then convert each leaf node into a sentence via a natural language template in some order, forming the text corresponding to the body of the problem. The default order follows the proof tree’s canonical ordering, but other orders are possible as well. The sentences are sampled uniformly at random from a predefined set of templates; some example templates are given in Table 1. The question is converted from the logical form at the root of the proof tree and is sampled from a set of interrogative sentences.
@@
In the final step, we generate the CoT annotation for the problem. We first obtain a sequential ordering of the CoT proof steps by visiting the nodes in the proof tree with a post-order traversal. [^9] Note that the choice of traversing the nodes in post order corresponds to applying each inference rule as soon as all of its premises are available. In addition, post-order traversal visits the axioms in canonical ordering. Alternative orders of tree traversal would violate at least one of these constraints. We then convert each proof step into a sentence via a natural language template. The templates used for the axioms are identical to those used in the body of the problem text; that is, the proof steps corresponding to axioms simply repeat the relevant part of the problem text. For the other proof steps, we use templates designed to explain the computation step that is performed. Fig. 1 illustrates the mapping between the nodes in the proof tree and the sentences in the CoT annotation. For instance, the third sentence in Fig. 1 explains the conclusion cont(Lucy, 27, apple), which is deduced from the two leftmost nodes in the tree: “So Lucy has $17+10=27$ apples”.
@@ 4.2 MathGAP: Mathematical Generalization on Arithmetic Proofs
With our generation method we can generate new synthetic problems where we control for the structure of the proof and its complexity. This mitigates the risk that the problems are contaminated. MathGAP can be used for various forms of systematic studies on arithmetic reasoning; here, we focus on in-context easy-to-hard OOD generalization. Specifically, we select a family of problems of interest (e.g., linear problems under comp inference rules) and generate a train-test split of problems in which problems in the test set are more complex than those in the training set (e.g., requiring deeper proofs).
The generation method can be flexibly adapted. One can add new logical forms and inference rules, restrict the choice of existing ones during sampling, or create logical forms for irrelevant sentences (Shi et al., 2023). One can also vary the text templates for the logical forms and inference rules, the vocabularies for agents, entities, attributes, and units, and the range of numbers from which the quantities are sampled. In addition, one may increase lexical and syntactic diversity by using an LLM to paraphrase the templated texts, e.g., through a procedure like Mishra et al.’s (2024). We do not perform such paraphrasing for the data in our experiments since it could introduce bias (Boguraev et al., 2024; Opedal et al., 2024) and produce problems that are unfaithful to their formal specifications.
5 Generalizing to Complex Proofs
We apply MathGAP to study how LLMs generalize from proof demonstrations of simple problems in context to solve problems of higher complexity at inference. We test OOD generalization through several sets of experiments: depth generalization for linear problems (§ 5.1), width generalization for linear problems (§ 5.2), depth generalization for nonlinear problems (§ 5.3), and generalization to permutations of the axioms (§ 5.4). Our public code repository provides further problem variations.
@@ 5 Generalizing to Complex Proofs
For each set of experiments, the general experimental setup is as follows: We generate multiple test sets of different degrees of complexity with 400 problems in each. We then generate model predictions for the problems in these test sets under four different prompts: (i) A zero-shot baseline prompt, which contains only the test problem; (ii) A prompt of primitive examples, in which all problems contain only one proof step of the same inference rule that is used in the test problem; (iii) A prompt of examples within a range of complexities, all of which are simpler than the test problem; (iv) An in-distribution baseline prompt, containing examples that are of the same complexity as the test problem.
Note that prompts (i) through (iii) evaluate in-context OOD generalization, since the test examples are from a distribution that is different from that of the in-context examples. For prompts (ii) through (iv) we generate a new set of in-context examples for every test problem. This is done in order to avoid bias in the results towards any particular set of examples. We set the number of examples to 12, except for the experiments on nonlinear problems for which we use 5—see App. C for the rationale. Importantly, the in-context problems are provided together with their CoT solution annotations, as generated through the MathGAP pipeline. For prompt (iii), we make sure that there is at least one problem of every complexity in the range. This prompt can be viewed as a form of in-context curriculum learning (Bengio et al., 2009), [^10] Related approaches have been explored in previous work (Li et al., 2022; Liu et al., 2024b). except that we randomize the ordering of the examples rather than arranging them in increasing order of complexity. This is done since in-context orderings can have a large effect on performance (Lu et al., 2022), and we do not want our results to overfit to any particular ordering. More details on the data and prompt are given in App. C.
Responses are generated using greedy decoding and a maximum context length of 4,096 tokens. Model predictions are obtained by extracting the last number occurring in the model output. We report answer accuracies and compute 95% confidence intervals using bootstrap sampling (Efron, 1992). We evaluate the following models on all test sets: Mixtral-8x7B (Jiang et al., 2024a), Llama3 with 8B and 70B parameters (Llama Team, 2024), GPT-3.5 Turbo and GPT-4o (OpenAI, 2024).
Figure 2: Answer accuracies for generalization to increasing depth and width for linear problems across models and in-context distributions. Depth is increased using inference rules involving comp and transfer . Width is increased using partwhole. See App. E for example problems.
@@ 5.1 Linear Depth Generalization
We consider linear problems using both comp and transfer . The five test sets consist of problems with depths between $6{-}10$. As such, the examples provided in the range of setting (iii) described above have depths between $1{-}5$. We give an example of a problem with depth $6$ in § E.1.
@@ Results.
The results are illustrated in the top and middle rows of Fig. 2. As expected, we observe a decreasing trend in performance as depth increases, for all in-context distributions. Curiously, GPT-3.5-Turbo is the only model for which adding in-context examples consistently leads to improvements. For the majority of cases, performance tends to be larger when providing a range of examples of varying complexity as compared to only primitive examples. The ranged OOD cases often show comparable performance to that of the in-distribution context, suggesting that these models benefit from demonstrations of problems with varying complexity. Llama3-8B is an outlier in that its solving accuracy for in-distribution context is (substantially) lower as compared to other prompts. When inspecting the outputs we see that the model often fails to answer the inference question, recognizing only that it has been provided a sequence of in-context examples. This could potentially be a consequence of how Llama3 is trained, using smaller context windows for the initial stages of pretraining (Llama Team, 2024). This appears to be remedied by scale since Llama3 70B shows near perfect performance.
@@ 5.2 Linear Width Generalization
We test generalization to larger proof width using partwhole problems with a fixed depth of $1$. The five test sets have widths between $7{-}11$ and the problems in setting (iii) described above have widths between $2{-}6$. Note that this matches the width for the problems in § 5.1; for those problems the width was equal to the depth plus one. § E.2 gives an example problem.
@@ Results.
The results are shown in the bottom row of Fig. 2. Apart from the GPT models, the performance is generally lower as compared to deep problems with the same number of sentences (§ 5.1), often with a sharper decreasing trend. Such is the case even for Llama3 70B, which performed almost perfectly in the linear depth setting. This suggests that it is more difficult to generalize in terms of width than in terms of depth, which is somewhat surprising considering that these problems only have one proof step. For GPT-3.5 and Mixtral-8x7B, we observe that providing in-context examples improves accuracy over zero-shot in most cases. As for linear depth generalization, it seems that having a range of OOD examples of varying complexity is superior to only including primitive examples. For the Llama models, adding OOD in-context examples often leads to lower performance over zero-shot.
Figure 3: Answer accuracies for generalization to increasing depth and width for nonlinear problems across models and in-context distributions. Depth is increased using inference rules involving comp and comp-eq . The in-distribution contexts were too large to fit in context for the Llama3 models and GPT-3.5 Turbo at some of the higher depths. App. F shows results on o1-preview and DeepSeek-R1.